OpenStudy (anonymous):

The sum of three consecutive integers is 267. What is the largest integer?

OpenStudy (anonymous):

n + n + 1 + n + 2 = 267. 3n + 3 = 267 n = ?

OpenStudy (anonymous):

you can call the integer \(n\) then the next one is \(n+1\) and the one after that is \(n+2\) add to get \[n+n+1+n+2=267\]

OpenStudy (anonymous):

Largest will be n+2.

OpenStudy (anonymous):

88+2=90 largest

OpenStudy (anonymous):

Sounds good.

OpenStudy (anonymous):

Although, we were not testing YOU to see if you can figure out from where we left off.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

wait so what was it.......

OpenStudy (anonymous):

lol let one integer x so consecutive will be one bigger than x= x+1 and other consecutive will be one bigger than x+1=x+2 now sum of these 3 = 267 x+x+1+x+2=267 3x+3=267 3x=264 x=88 now we know that the biggest one is x+2 right? plus in the value of x 88+2=90 which is biggest

OpenStudy (anonymous):

plug in*

OpenStudy (anonymous):

so its 90? or is it that because its supposed to be a numerical number